If Courtney were three times as old as she was six years ago, she will be fifty- four less than five times her current age. How old is Courtney?
Answer by richwmiller(17219)

3(a-6)=5a-54
3a-18=5a-54
54-18=2a
27-9=a
18=a
check
3*12=5*18-54
36=90-54
36=36
ok